Please Help:
Find the present value of the future amount. Assume 365 days in a year. Round to the nearest cent.
$500 for 298 days; money earns 6%
A) $471.70
B) $476.73
C) $23.35
D) $476.65

Answer by rfer(16322)

first find the interest rate for 298 days.
.06/365=.000164383*298=.0489863
500/1.0489863=476.65 (D)